; Let at be a linear transformation as follows : T{x1,x2,x3,x4,x5} = {{x1-x3+2x2x5},{x2-x3+2x5},{x1+x2-2x3+x4+2x5},{2x2-2x3+x4+2x5}]

a.) find the standard matrix representation A of T

b.) find the basis of Col(A)

c.) find a basis of Null(A)

d.) is T 1-1? Is T onto?
Answer #1

T(x1, 22, 23, 24, 25) = {(x,- x₂ + 2x2 + xg), CX2 X3 +285), (x, +82-2X2 + x4 +2xg), (2x2 - 2x3 +24+225)} find the standard maA -1 OL -1 o 2 Tooror T T 2 O -2 L - 2 t оо 2 2 o find the basis of collas convert A into echelone from R3 R₃-RI T 2 1 - 1 0a basix of Null (A) (co find Parc = $(x,x2, 11317*83 | TCX, 42,43,44) = (0,0,0,0)) (x, 7X2-2X ₂ +X 4+28), (282-283 +24+25) >(x+222-x₂+x=0 X2-23 + 2x5=0 - 2x 3+x4+x=0 xq-2x=0 X4=28 - 2X3 + 2Xg+3x70 x, tts -Sxg+x=0 x + (455) XS=0 xs 2 12-+235-0 x 2 Nulbecaus Kealt) is not (d) T is not one-one equal to jero. ie. Hull space have Monjeno tem etes Vector but I is onto because Ra